Give the equation for the circle with center C(3,-2) and radius 4. Thanks!
Answer by AnlytcPhil(1810)

Give the equation for the circle with center
 C(3,-2) and radius 4. Thanks!

You just have to learn the equation

(x-h)² + (y-k)² = r²

Where the center = (h,k) = (3,-2)

Substitute 3 for h and -2 for k and 4 for the radius r.

Then you have

(x-3)² + (y-(-2))² = 4²

You can simplify

(x-3)² + (y+2)² = 16

And you can multiply it out if you like

(x-3)(x-3) + (y+2)(y+2) = 16

x² - 3x + 9 + y² + 4y + 4 = 16

x² + y² - 3x + 4y + 13 = 16

x² + y² - 3x + 4y - 3 = 0

Edwin